how do you find the length of the altitude of a triangle when given the hypotenuse. Used for 30-6-90 degree triangle. and radical form. thanks
Answer by edjones(8007)

All 30, 60, 90 triangles are similar in this ratio: base=1, altitude=sqrt(3), hypotenuse=2.
Example: base=5 then altitude=5sqrt(3) and hypotenuse=5*2=10
.
Ed